OpenStudy (anonymous):

Exponential decay finding r 48=64(1-r)^10

OpenStudy (anonymous):

\[(1-r)^{10} = \frac{48}{64}\]

OpenStudy (anonymous):

\[(1-r) = \sqrt[10]{\frac{48}{64}}\]

OpenStudy (anonymous):

\[(1-r) = +- \sqrt[10]{\frac{48}{64}}\]

OpenStudy (anonymous):

then rearrange, its easy